Solve each system by substitution. Check your solution.
19.
a-1.2b=-3
0.2b+0.6a=12

Found 2 solutions by checkley77, richwmiller:
Answer by checkley77(12844)

a-1.2b=-3 or a=1.2b-3
0.2b+0.6a=12
.2b+.6(1.2b-3)=12
.2b+.72b-1.8=12
.92b=12+1.8
.92b=13.8
b=13.8/.92
b=15 ans.
a=1.2*15-3
a=18-3
a=15 ans.
Proof:
.2*15+.6*15=12
3+9=12
12=12

Answer by richwmiller(17219)

a-1.2b=-3
0.2b+0.6a=12
we have two linear equation which form lines on a graph with two variables
we can then solve for the two variables by elimination or substitution.
we will substitute
a-1.2b=-3
add 1.2b to both sides
a=1.2b-3
Substitute 1.2b-3 for a in the second equation.
0.2b+0.6(1.2b-3)=12
expand and solve for b
.2b+.6*1.2b-.6*3=12
add .6*3 to both sides
.2b+.6*1.2b=.6*3+12